OpenStudy (anonymous):

Find an equation of the line containing the given pair of points (-3,0) and (0,4) y=?

OpenStudy (anonymous):

Should have got this by now, get the slope and then use one of the points to get the line.

OpenStudy (anonymous):

Ok, tell me if I am right...... 4-0=4, and 0- -3=-3, so 4 / -3 is 1.333

OpenStudy (anonymous):

0--3 is 3 (negative minus or multiply negative is positive)

OpenStudy (anonymous):

but the solution is still 1.33333333

OpenStudy (anonymous):

for the slope.

OpenStudy (anonymous):

Usually it is better to leave it as a fraction

OpenStudy (anonymous):

1/3

OpenStudy (anonymous):

4/3

OpenStudy (anonymous):

I have to put this in slope-intercept form

OpenStudy (anonymous):

You have got the slope, do u know how to get y= using this and one of the points?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

OK, let's use (0,4) then y-4 = 4/3(x-0) -> y = 4/3 x +4